As the price of ticket rises from $200 to $300, what is the price elasticity of demand for

(i) Business travelers

(ii) Tourists

(Use the midpoint method in your calculations.)


 price

Quantity demanded (business travelers)

Quantity demanded (tourists)

$150

2100

1000

200

2000

800

250

1900

600

300

1800

400

 


 


1
Expert's answer
2021-01-29T20:00:23-0500

(i) Let's calculate the price elasticity of demand for business travelers:


"E_d=\dfrac{\dfrac{Q_2-Q_1}{0.5(Q_2+Q_1)}}{\dfrac{P_2-P_1}{0.5(P_2+P_1)}},""E_d=\dfrac{\dfrac{1800-2000}{0.5(1800+2000)}}{\dfrac{300-200}{0.5(300+200)}}=-0.26"

(ii) Let's calculate the price elasticity of demand for tourists:


"E_d=\dfrac{\dfrac{Q_2-Q_1}{0.5(Q_2+Q_1)}}{\dfrac{P_2-P_1}{0.5(P_2+P_1)}},""E_d=\dfrac{\dfrac{400-800}{0.5(800+400)}}{\dfrac{300-200}{0.5(300+200)}}=-1.66"